Shift report 20030121 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Jan 21 2003 - 00:17:41 EST

  • Next message: Ian Bearden: "HPSS database problem?"
    ------------------------------------------------------------
    Date: Tue Jan 21 00:17:40 2003, Shift: 16:00-24:00
    Supervisor: AM, with: 
    Log book#13 pages 107 - 108
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam for the most part of the shift, but backgrnd got high
    pretty often, what caused several DAQ crashes.
    2. Runs 7518, 7519, 7520, 7521, 7522 have been recorded
    3. DAQ crahsed several times during the shift, mostly when
    backgrnd got too heigh. No other detector problems were observed.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Jan 21 2003 - 00:21:26 EST